### v1

# Accelerating Large Language Model Pretraining via LFR Pedagogy: Learn, Focus, and Review

Abstract
Large Language Model (LLM) pretraining traditionally relies on autoregressive language modeling on randomly sampled data blocks from web-scale datasets. We take inspiration from human learning techniques like spaced repetition to hypothesize that random data sampling for LLMs leads to high training cost and low quality models which tend to forget data. In order to effectively commit web-scale information to long-term memory, we propose the LFR (Learn, Focus, and Review) pedagogy, a new dynamic training paradigm which focuses and repeatedly reviews complex data blocks at systematic intervals based on the model’s learning pace and progress. LFR records the model perplexities for different data blocks and frequently revisits blocks with higher perplexity which are more likely to be forgotten. We pretrain the GPT-2 models (124M - 1.5B) from scratch on the OpenWebText dataset using LFR. We test on downstream tasks from the language modeling, question answering, translation, and problem solving domains to achieve consistently lower perplexity and higher accuracy than the baseline OpenAI models, while obtaining a 20x pretraining speed-up.
Accelerating Large Language Model Pretraining via LFR Pedagogy: Learn, Focus, and Review
Neha Prakriya UCLA nehaprakriya@cs.ucla.edu                        Jui-Nan Yen UCLA juinanyen@cs.ucla.edu
Cho-Jui Hsieh Google and UCLA chohsieh@cs.ucla.edu                        Jason Cong UCLA cong@cs.ucla.edu
1 Introduction
LLMs have achieved remarkable success in understanding and generating human language. This success is driven by the ever-increasing model parameter sizes which require web-scale training datasets like OpenWebText Radford et al. (2019); ope , CommonCrawl Penedo et al. (2023); Raffel et al. (2023), and Pile Gao et al. (2020), leading to unsustainable training costs. For example, pretraining the GPT-4 model OpenAI et al. (2024) was estimated to have cost around \$100M USD over a period of 3-4 months using 25k NVIDIA A100 GPUs gpt . As such, a key challenge for unlocking the next generation of language models is to significantly reduce training costs while retaining or improving downstream task performance.
Popular LLMs, like the GPT series Radford et al. (2019); Brown et al. (2020); OpenAI et al. (2024), Llama series Touvron et al. (2023, 2023); lla (2024), Mistral Jiang et al. (2023), and others, were trained utilizing autoregressive/causal techniques, where the goal of training is to generate a plausible next token based on a randomly sampled input data block from the training corpus. However, data scraped from the web is highly redundant and presents varying complexity. For example, Table 1 shows two randomly selected example texts from the OpenWebText Radford et al. (2019); ope dataset. If a human is asked to learn these two texts, they might spend more time learning the details in the first block of text than those in the second. The first block contains more complex information with facts and dates, while the second block contains anecdotal information, which can be understood faster. They might read the first example more times than the second. However, autoregressive language modeling does not mimic such behavior.
The slave rebellion of the early 70s BCE led by Spartacus, the Thracian gladiator, would present Crassus, made praetor in 73 BCE, with an opportunity to flex his military muscle and gain further prestige with the Roman people. The slave army, numbering between 70,000 and 120,000, was a serious threat and had already defeated two separate Roman armies and two consuls. Now they were ravaging the southern Italian countryside and it was Crassus who was entrusted with finally removing this thorn from Rome’s heel.
Sata, who is older than Japan’s nearly 60-year-old civilian nuclear industry, remembers a time without nuclear power. Families were allowed just a few lightbulbs in the 1940s, because the electrical system was still in its infancy. “There was a TEPCO office in the neighborhood,” she recalls, and when a bulb burned out, “we had to bring it there” to trade it for a new one. The advent of nuclear power meant that the Japanese could consume as much electricity as they wanted.
Table 1: Examples of Text Blocks in the OpenWebText Dataset
Inspired by these observations on different learning methods used by humans and LLMs, our work has the following contributions:
  1. 1.
    Profile LLM pretraining to observe multiple descent behaviour for 25% of the training tokens that are forgotten multiple times.
  2. 2.
    Develop a Learn-Focus-Review (LFR) training pipeline that uses perplexity to gauge the LLM’s learning pace, focusing on complex data blocks while regularly reviewing all data blocks to prevent forgetting.
  3. 3.
    Pretrain GPT-2 models (124M-1.5B) on the OpenWebText dataset with 4 AMD MI210 (345M-1.5B) and 4 AMD MI100 GPUs (124M), where the 1.5B model training took less than 20 days, and evaluate them on six downstream tasks covering language modeling, question answering, translation, and problem solving. In fact, we believe that we are the first to train the GPT-2 models on a cluster of AMD GPUs, demonstrating an alternative GPU platform for LLM training.
  4. 4.
    LFR results in significantly lower perplexity and higher accuracy compared to baseline OpenAI models, achieving these improvements with 20x fewer training iterations.
  5. 5.
    Observe that LLMs first learn conversational and anecdotal data, before being able to retain factual and instructional information in long-term memory.
  6. 6.
    Demonstrate that text importance varies with training time and model size, driving the need for dynamic data selection methods like LFR.
In the following sections, we examine prior works on efficient LLM pretraining before diving deeper into our proposed training strategies and design decisions.
2 Related Work
Prior works on efficient pretraining of LLMs using data selection have primarily focused on using distance metrics and clustering techniques. Tirumala et al. (2023) proposes D4, which deduplicates and selects cluster centers in the embedding space generated by pretrained models. SemDeDup Abbas et al. (2023) prunes semantic duplicates using pretrained models. It can successfully prune 50% of the training data with minimal performance loss. MiniPile Kaddour (2023); min uses the pretrained E5-Large Wang et al. (2024) model to embed documents in the Pile dataset and clusters them to select a smaller pretraining corpus of $\sim$6GB. DSIR Xie et al. (2023) proposes selecting subsets from large unlabeled datasets through importance resampling to match the distribution of the target dataset. However, considering the high cost of training, it is unsustainable to sample a new subset and pretrain the LLM from scratch for every new downstream task. While these works address the common redundancy issue present in LLM pretraining corpuses, they face three major drawbacks. First, they require pretrained models for calculating the distance metric on embeddings. This step must ensure that the embeddings found by the pretrained model are similar to those of the target model to ensure the transferability of sample importance. Second, clustering-based methods do not scale to encompass the diversity of data in web-scale datasets. This can lead to a loss in downstream task accuracy (see Section 5). Third, they are slow and can take as long as 2-3 hours to generate 6GB subsets of the OpenWebText dataset (38GB) as shown in Section 5.
Marion et al. (2023) evaluates data selection by pretrained models used to obtain perplexity, Error L2-Norm (EL2N) Paul et al. (2021), and memorization rankings Biderman et al. (2023). They find that perplexity is a good metric to characterize a model’s understanding of a data sample and that retaining data samples with middle perplexity values can outperform training on the entire dataset. While this finding inspires our work, their method relies on pretrained model checkpoints to evaluate data sample importance. This limits the scalability.
3 Problem Formulation and Profiling
3.1 LLM Pretraining Objective
Given an LLM model parameterized by weights $\theta$ and a web-scale dataset $D$, we first tokenize all the documents in the dataset and obtain context-length sized sequences of tokens $s_{i}$ such that the training corpus becomes $D=\{s_{1},s_{2},s_{3},...s_{n}\}$. For the OpenWebText dataset used in this paper, the context-length is 1024 and there are a total of 9B training tokens. Given one such sequence of tokens $s_{i}=\{x_{1},x_{2},...x_{n}\}$, the training objective is to autoregressively predict the next $M$ tokens:
$p_{\theta}(y\mid x)=\prod_{i=1}^{M}p_{\theta}(y_{i}\mid y_{1:i-1},x).$ (1)
The batches of sequences ($s_{i}$) used as input to prompt the LLM (shown in Eq. 1) and backpropagate for each step are randomly sampled from the training corpus.
3.2 Observations from Training on Randomly Sampled Data
In order to better understand the drawbacks of this traditional training technique, we profile the pretraining process for the 345M parameter GPT-2 model. The training hyperparameters are provided in Appendix A.1. Similarly to Marion et al. (2023), we use perplexity as a metric to monitor the training progress. Given a token sequence $s_{i}=\{x_{1},x_{2},...,x_{n}\}$ from the dataset $D$, perplexity is computed as:
$PPL(s_{i})=\exp\left(\frac{1}{|s_{i}|}\sum_{x_{j}\in s_{i}}NLL(x_{j})\right),$ (2)
where $NLL(x_{j})$ is the negative log likelihood of token $x_{j}$ computed as follows:
$NLL(x_{j})=-\log P(x_{j}\mid x_{<j};\theta).$ (3)
Using this metric, models exhibiting lower perplexities are considered better since this indicates a high probability of selecting text closest to the raw dataset.
To generate profiling information, we record the data block ID (each data block consists of 1024 tokens) and model perplexity each time a block is randomly sampled from the training corpus across all training iterations. At the end of the training process, we have a record of the perplexities over time for each data block. Since we train the GPT-2 345M model for 8 epochs, each data block has at least 8 perplexity values captured at different training iterations. From this record, we classify each data block as one of the following:
  1. 1.
    Learned: recorded perplexities monotonically decrease.
  2. 2.
    Unlearned: recorded perplexities monotonically increase.
  3. 3.
    Forgotten: recorded perplexities first increase and then decrease. Such an upward and downward trend may repeat any number of times during training.
We observe that 25% of data blocks are forgotten at least once during training and that of the data blocks that are forgotten, 82% are forgotten multiple times during training, i.e., they display multiple descent behavior (Figure 1). Xia et al. (2022) reported a double-descent behavior for the OPT models Zhang et al. (2022), and our above experiment further demonstrates that the “forgotten” can happen multiple times in LLM pretraining.
Figure 1: Perplexities of data samples being forgotton by the GPT2-345M model.
4 LFR Training Methodology
Based on our profiling observations in Section 3.2 we propose to replace traditional autoregressive language modeling methods with Spaced Repetition Tabibian et al. (2019). Spaced Repetition is an evidence-based learning method proven to improve information retention and learning pace in humans Smolen et al. (2016). In this technique, challenging pieces of information are reviewed more often, at regular intervals, and easier pieces of information are presented to the learner less often. We pretrain our models with a combination of the following three steps:
  1. 1.
    Learn: Initially, we present the model with the entire dataset and train on randomly selected data blocks for $p_{1}$ epochs, as normally seen in the traditional approach. $p_{1}$ can be configured by the user based on the available compute budget, model, and dataset. During the $p_{1}$ epochs, we also record the perplexities for all data blocks. The next 2 phases can be repeated $reps$ number of times based on the available compute budget.
  2. 2.
    Focus: From the collected perplexities from the previous phase, we discard $s_{1}$of the data blocks corresponding with the lowest perplexities. In doing so, we provide a mechanism for shifting the model’s focus towards learning data blocks that were determined to be difficult. In this phase of training, we restrict the random sampling of data points to this reduced subset for $p_{2}$ epochs. $s_{1}$ and $p_{2}$ are user-controlled hyperparameters.
  3. 3.
    Review: Next, we reintroduce all of the removed data blocks and train the model on the entire corpus for $p_{3}$ epochs (reconfigurable). This ensures that we allow the model to review and revisit data blocks which it may have forgotten. We record the perplexities for all data blocks again in this training phase.
Our algorithm is detailed in Algorithm 1. For pretraining the GPT-2 models, we choose the following combination of the aforementioned steps:
  1. 1.
    Phase 1: Learn for 1 epoch ($p_{1}=1$).
  2. 2.
    Phase 2: Focus on 50% of the data for 1 epoch ($s_{1}=50,p_{2}=1$).
  3. 3.
    Phase 3: Review the entire dataset for another epoch ($p_{3}=1$).
  4. 4.
    Phase 4: Focus on 30% of the data for 5 epochs ($reps=2,s_{2}=70,p_{4}=1$).
We choose this configuration based on our limited pretraining budget and the profiling in Section 3.2 which revealed that $\sim$25of data samples are forgotten at least once during training. From our observations in Figure 1, we notice that the data samples that are forgotten are forgotten multiple times. On average, they had to be presented to the model 4 times for them to be learned. Therefore, we use the first three training phases to identify these samples, and then spend 5 epochs focusing on them. To be on the conservative side, we focus on 30% of the samples in the last 5 epochs, to ensure that the model retains these difficult samples in its long-term memory. Furthermore, we ensure that all data samples are reviewed through Phases 1 and 3 of our approach. Finally, we test other values for $p_{1}$, $s_{1}$, $p_{2}$, $p_{3}$, and $reps$ and evaluate the sensitivity of LFR to these hyperparameters in Section 5.9.
Algorithm 1 LFR Training Methodology
Training dataset $D$, model $M$ with initial parameters $\theta_{0}$, hyperparameters $p_{1}$, $s_{1}$, $p_{2}$, $p_{3}$, and $reps$.
Minimization of Equation 3.
$PPLs,\theta_{p_{1}}\leftarrow\textbf{Learn}(\theta_{0},D,p_{1})$
for $r=1,2,\ldots,reps$ do
     $Sort(PPLs,D),S_{1}\leftarrow(100-s_{1})\%D$
     $\theta_{p_{2}}\leftarrow\textbf{Focus}(\theta_{p_{1}},S_{1},p_{2})$
     $PPLs,\theta_{p_{3}}\leftarrow\textbf{Review}(\theta_{p_{2}},D,p_{3})$
end for
$Return\ \theta$
Our training strategy is simple, intuitive and human-like. It gives the model an opportunity to learn from all of the data, prioritize and relearn forgotten data points, and repeatedly review harder data samples more frequently than they would have been using random sampling. While the clustering-based techniques Tirumala et al. (2023); Abbas et al. (2023); Kaddour (2023) presented in Section 2 allow for accelerated training, they are not designed to suit the multi-descent training dynamics observed in Section 4 and require pretrained model embeddings to calculate distance metrics. By dropping data points with high similarity (within the same cluster), the model is not given a chance to repeatedly learn harder data points and observe the multi-descent behaviour in Figure 1. Furthermore, prior methods including perplexity-based pruning methods Marion et al. (2023) are static. Sections 5.8 and A.2 characterize the data blocks found easy and hard by the LLM, and demonstrate why static, clustering-based data selection methods achieve poor downstream task performance. Lastly, our approach does not require any pretrained models to obtain embeddings. Our focused training strategy allows the model to absorb harder information (data blocks with higher perplexity) faster, by presenting them more number of times. We evaluate the performance of LFR on six downstream tasks in Section 5.
5 Evaluation
In this section, we present a comprehensive evaluation of our proposed data selection method, LFR. We pretrain GPT-2 models Radford et al. (2019) of various sizes (124M, 345M, 774M, 1.5B) on the OpenWebText dataset ope from scratch using 4 AMD MI210 GPUs (345M-1.5B) and 4 AMD MI100 GPUs (124M). Our pretraining experiments utilize a fully distributed data parallel (DDP) approach.
We also evaluate our pretrained models on downstream tasks from different domains in Sections 5.3, 5.4, and 5.5 by finetuning it with the Huggingface transformers library Wolf et al. (2020). Section 5.8 presents an analysis on the data blocks dropped and retained by LFR in different phases of training across the different model sizes. Finally, we conduct a sensitivity study in Section 5.9, where we test LFR with different hyperparameters.
5.1 Models
The GPT-2 models implement a decoder-only transformer architecture introduced in Radford et al. (2019). The model configurations are shown in Table 6 in Section A. The models were originally pretrained on the closed-source WebText dataset Radford et al. (2019) for 800k iterations. Researchers developed the OpenWebText dataset as an open-source clone of the WebText dataset by following the same data scraping method described by OpenAI in their GPT-2 release Radford et al. (2019). LFR utilizes the OpenWebText dataset for pretraining for a total of 40k iterations, i.e., 20x fewer iterations than the OpenAI release. We build upon the GPT-2 training implementation provided by Nan and train our models with the four phases described in Section 4.
Model WikiText-2 WikiText-103 LAMBADA 1BW
124M-OpenAI (800k iters) 22.1 31.58 18 39.18
124M-RS (40k iters) 23.32 23.42 17.71 39.49
124M-LFR (40k iters) 19.81 22.49 16.61 32.27
345M-OpenAI (800k iters) 19.82 22.05 14.26 29.95
345M-RS (40k iters) 21.11 21.8 14.84 30.66
345M-LFR (40k iters) 16.31 17.48 13.7 25.52
774M-OpenAI (800k iters) 15.93 18.53 13.74 26.52
774M-RS (40k iters) 16.71 18.89 14.10 28.56
774M-LFR (40k iters) 15.11 14.58 12.51 23.83
1.5B-OpenAI (800k iters) 13.80 16.59 12.15 23.87
1.5B-LFR (40k iters) 13.10 14.37 11.23 22.09
Table 2: $PPL$ results for language modeling datasets across model sizes. Here, $N$-OpenAI refers to the OpenAI baseline (trained for 800k iterations), $N$-RS refers to the random sampling baseline (trained for 40k iterations), and $N$-LFR refers to our proposed training pedagogy (trained for 40k iterations), where $N$ is the number of model parameters.
5.2 Baselines
We evaluate the models pretrained using LFR with a comprehensive set of prior works and industry-standard checkpoints. They include:
  1. 1.
    OpenAI checkpoints: we compare our models pretrained through LFR for 40k iterations with the same GPT-2 architectures pretrained by OpenAI for 800k iterations. We use the same batch size as OpenAI (512) as shown in Section A.1. Due to insufficient GPU memory, we adjust the batch size by using a higher gradient accumulation. The checkpoints are sourced from Huggingface [^1] https://huggingface.co/openai-community.
  2. 2.
    Random Sampling: while the previous baseline ensures that we compare with industry-standard models, we also develop and compare LFR against the same GPT-2 architectures pretrained for 40k iterations on the OpenWebText dataset using random sampling. This baseline enables LFR to produce higher quality models than those obtained through traditional autoregressive modeling when using the exact same dataset and iteration count. Note that we are only able to produce and pretrain this baseline for the 124M, 345M, and 774M models due to limited compute resources.
  3. 3.
    Prior works: We pretrain the GPT-2 124M parameter models on the dataset generated through DSIR Xie et al. (2023) and MiniPile Kaddour (2023). More details of these baselines can be found in Section 5.7.
5.3 Language Modeling
To evaluate the performance of a retrained language model, we finetune on a variety of datasets and evaluate the performance in terms of $PPL$. Specifically, we consider WikiText-2 wik , WikiText-103 wik , LAMBADA Paperno et al. (2016), and One Billion Word Benchmark (1BW) Chelba et al. (2014) and show the results of our method and baselines in Table 2.
Compared with the OpenAI baseline, we observe large improvements with a decrease in $PPL$ ranging from 0.82-3.51 (5.3%-17.7%) for the WikiText-2 dataset across model sizes. The improvements are more significant for the larger WikiText-103 dataset. We observe a $PPL$ reduction by 3.58-9.09 (20.7%-28.8%) across model sizes. In the case of the LAMBADA dataset which measures the LLM’s ability to capture long-range dependencies in text, LFR’s focused learning approach outperforms the OpenAI checkpoints by lowering perplexity by 0.56-1.39 (3.9%-8.9%). Finally, for the 1BW benchmark, which measures progress in statistical language modeling, LFR outperforms the OpenAI models by a large reduction in $PPL$ ranging from 2.69-6.91 (10.1%-17.6%).
Based on the results in Table 2, we find that the $PPL$ reduction obtained by LFR increase as the dataset size increases (from WikiText-2 to 1BW). Also, smaller models show a larger $PPL$ reduction by using LFR than larger models. On average, using our approach, perplexity was reduced by 4.92, 3.26, 2.17, and 1.40 for the GPT-2 124M, 345M, 774M, and 1.5B models, respectively.
5.4 Translation
The WMT-14 English-French dataset Artetxe et al. (2018); wmt consists of pairs of sentences in English and French. While, non-English content is filtered out in the OpenWebText dataset (similarly as in WebText), the GPT-2 1.5B parameter model pretrained through LFR can achieve a BLEU score of 11.8 on the WMT-14 English-French test set. In contrast, the GPT-2 1.5B OpenAI baseline achieves a lower BLEU score of 11.5. Note that the GPT-2 models are not designed for translation tasks, but can still outperform unsupervised machine translation approaches Artetxe et al. (2018).
5.5 Question-Answering
Question-answering is an important domain for LLMs and is a potential test for the factual information aggregated in the parameters during training. Similarly to the GPT-2 paper, we test our pretrained models on the Natural Questions dataset Kwiatkowski et al. (2019) by augmenting our model with a linear layer to compute the start and end logits of the model output. The GPT-2 1.5B parameter baseline OpenAI model achieves 4.1% accuracy on the questions while LFR achieves a slightly higher accuracy of 4.61%. Considering that model capacity plays a major role in improving accuracy for question-answering tasks, we believe that applying LFR’s training methodology to larger models will yield even better results.
5.6 World Knowledge and Problem Solving
We evaluated LFR on the MMLU benchmark Hendrycks et al. (2021) to test the model’s ability to solve multiple-choice questions across a wide range of 57 subjects, including mathematics, US history, computer science, law, and social sciences. We compared the GPT-2 1.5B model, trained with LFR, against the OpenAI baseline in a five-shot setting. This resulted in an average accuracy improvement of 0.2% across all 57 subjects. Table 3 displays the accuracy results.
Subject 1.5B-OpenAI 1.5B-LFR STEM 24.5% 26.1% Humanities 24.8% 27.2% Social Sciences 24% 23.8% Other (business, health, misc.) 27.8% 25.1% Average (across 57 subjects) 25.3% 25.5%
Table 3: Accuracy results for the MMLU benchmark for LFR and the OpenAI Baseline for the 1.5B parameter model.
5.7 Comparison with Prior Work
WikiText-2 WikiText-103 Lambada 1BW DSIR 22.6 25.29 17.34 35.33 MiniPile 150 322.1 73.61 459.2 LFR 19.81 22.49 16.61 32.27
Table 4: $PPL$ achieved by LFR and prior work on data selection for accelerated LLM training.
In this section, we use the 124M parameter model to compare LFR with DSIR Xie et al. (2023) and MiniPile Kaddour (2023), which propose importance resampling and cluster-based data sampling respectively. In the case of DSIR, we use the OpenWebText dataset as the raw dataset and set the target dataset as WikiText-103 for selecting a subset of 1M documents ($\sim$6GB). It took a total of 2.7 hours for the data selection process on 8 AMD EPYC 7V13 CPU cores operating at 2.45GHz with 64GB RAM. For the MiniPile baseline, we pretrain the GPT-2 124M model for 40k iterations with the dataset available on Huggingface (min, ). To ensure a fair comparison, we finetune each model using 4 AMD MI100 GPUs on the same downstream tasks used in Section 5.3. Table 4 displays the downstream task perplexity. LFR significantly outperforms prior work on data selection and allows the pretrained LLM to generalize better to new finetuning objectives, by focusing and revising different parts of the training corpus at regular intervals.
Figure 2: Cosine similarity heatmaps for dropped data blocks during phases 2 and 4 of pretraining for the GPT-2 124M (right) and 345M (left) models. The smaller model displays greater similarity in dropped data blocks over time (lighter color), indicating that it remained uncertain about similar data points than the larger model.
Figure 3: Cosine similarity heatmap for data blocks dropped during Phase 2 of GPT-2 124M and 345M pretraining shows moderate similarity, indicating different data points are considered easy by each model.
5.8 Analysis on Dropped and Retained Data Blocks
In this section, our goal is to characterize the data points retained and dropped during pretraining by LFR in Phases 2 and 4 (Sec. 4) across the training time and model size. Specifically, we aim to answer the following questions:
  1. 1.
    What types of data blocks are learned earlier in the training process compared to those learned later?
  2. 2.
    Are similar data blocks considered learned and dropped in Phases 2 and 4?
  3. 3.
    Are the dropped data blocks similar across model sizes?
  4. 4.
    Are the data blocks dropped similar to those retained at any given training phase?
To answer the first question, we printed out the texts dropped and retained at different training phases. Tables 12 and 14 show text blocks dropped in Phases 2 and 4 by the 345M and 124M parameter models, while Tables 13 and 15 show data blocks retained. By reading through the texts, we notice that the model first learned conversations and personal anecdotes, before being able to retain factual information. We provide a more detailed analysis of the learning process in Section A.2.
In order to answer questions 2-4, we recorded only the IDs of dropped data blocks during Phases 2 and 4 for both the GPT-2 124M and GPT-2 345M models, totaling 4 lists of dropped IDs. We then load the recorded data blocks and embed them into a higher dimensional space using the GPT-2 tokenizer. Considering that there is a total of 8.7M data blocks (9B tokens divided into blocks of 1024 tokens), we cluster the embeddings using $k$-means clustering with $k=270$ to reduce the analysis space and complexity. Finally, for each model, we compute the cosine similarity for all combinations of the embeddings of dropped data blocks across training phases and visualize them using a heatmap. These heatmaps plot the cosine similarity values (ranging between 0 and 1) such that lighter values (closer to 1) indicate higher similarity.
Figure 2 shows the similarity of dropped data blocks across the time scale (Phase 2 shown on the X-axis and Phase 4 shown on the Y-axis) for the 124M (left) and 345M (right) parameter models. We find that there is a higher similarity in the data points dropped by the 124M parameter model in Phases 2 and 4 than in the case of the 345M parameter model (mean, variance, and standard deviation are provided in Table 9). This behavior signals that the lower capacity of the 124M parameter model inhibits its learning process in Phase 3, such that it finds similar points confusing in Phases 2 and 4. In contrast, the 345M parameter model learns the data blocks it found confusing in Phase 2 by focusing on them, and moves on to learning new complex blocks by Phase 4.
We conduct a similar study in order to characterize the similarity in data blocks across model sizes. Figure 3 plots the cosine similarity heatmap for the data blocks dropped by the 124M parameter model (X-axis) and those dropped by the 345M parameter model (Y-axis) in Phase 2. The mean, variance, and standard deviations of the cosine similarity are 0.38, 0.15, and 0.023, respectively. This indicates that the data blocks found easy and dropped in Phase 2 by both models display a moderate level of similarity, but also differ significantly.
Finally, we observe the cosine similarity of data blocks dropped and retained during phase 4 for the 124M (left) and 345M (right) parameter models in Figure 4. The mean, standard deviation, and variance are detailed in Table 10. The smaller model displays greater similarity (lighter values in the heatmap) between the dropped and retained blocks than the larger model. We hypothesize that the larger model can perform reasonably well across similar data points, but struggles with very different complex blocks by the fourth training phase. In contrast, the smaller model does not display the same high-level of understanding (similar perplexity values) on related data blocks.
To summarize, data block importance varies across training time, and across model sizes. Therefore, static data selection techniques Tirumala et al. (2023); Abbas et al. (2023); Kaddour (2023); Xie et al. (2023) which select a fixed subset to train for the entire training duration for all model architectures do not adapt to the changing training dynamics of LLMs. Based on our analysis in Figure 3 and 2, different data blocks are found difficult by models of different capacities at different training instants, driving the need for dynamic data selection methods like LFR. We detail further analysis on the selected and discarded data blocks and demonstrate how models initially focus on learning conversational and anecdotal data, before proceeding to learn factual data in Appendix A.2.
5.9 Sensitivity Study
In this section, our goal is to understand the effects of more aggressive focus, revision, and learning strategies than the 4-phase training strategy presented in Section 4. Here, we vary the values of hyperparameters $p_{1},s_{1},p_{2},p_{3},$ and $reps$ and study the effects on the downstream task perplexity. Specifically, we aim to answer the following two questions:
  1. 1.
    What is the impact of removing Phase 3?
  2. 2.
    What is the impact of the degree of pruning in Phases 2 and 4?
To answer the first question, we pretrain a 124M parameter GPT-2 model without the reintroduction of data blocks in Phase 3, and use the reduced subset from Phase 2 for the rest of the training. Then, we finetune for downstream language modeling tasks similarly to in Section 5.3 and compared the perplexities using LFR in Table 5. This training strategy which removes Phase 3, is labeled as no-reintro. Next, to answer the second question, we pretrain a 124M parameter GPT-2 model using LFR but increase the degree of pruning in Phase 2 from 50% to 70%, i.e., reduce the training subset to 30% of the original size. This aggressive training strategy is labeled as aggr-2.
We observe that both aggressive training strategies do not work as well as the original method. However, we continue to explore more automated ways of deciding the training schedule for different model families as part of our future work.
Model WikiText-2 WikiText-103 LAMBADA 1BW no-reintro 23.24 25.76 17.27 36.02 aggr-2 23.91 27.00 21.11 38.62 LFR 19.81 22.49 16.61 32.27
Table 5: Downstream task perplexities with more aggressive training strategies.
6 Conclusion
We present LFR, a new training paradigm which proposes a mechanism for models to learn, focus, revisit and review complex data during training. Our method is inspired by the effective learning methods used by humans. LFR achieves consistently lower perplexity and higher accuracy on downstream task performance over traditional autoregressive modeling-based methods while using 20x fewer training iterations. We also demonstrate that LLMs first learn anecdotal and conversational data, before being able to retain factual information.
7 Acknowledgements
This work is partially supported by the PRISM Center under the JUMP 2.0 Program, AMD HACC Program, and NSF 2048280, 2331966, 2325121, 2244760. The authors would like to thank AMD for their generous support provided through the AMD AI & HPC Fund Research Award.
8 Limitations and Ethical Considerations
LFR presents the following directions for future work:
  1. 1.
    LFR is evaluated on models up to 1.5B parameters using the OpenWebText dataset, constrained by our compute resources. We encourage researchers to validate such focused learning approaches for different model families, and domains.
  2. 2.
    The sensitivity study in Section 5.9 reveals that the hyperparameters selected in Section 4 have a large impact on the performance of the trained model. Due to our limited compute budget, we are unable to present more comprehensive hyperparameter tuning experiments than those presented in Section 5.9.
Appendix A Appendix
A.1 Experiment Details
Datasets The datasets used for our experiments in Section 5 are detailed below:
  1. 1.
    WikiText wik : the WikiText language modeling dataset consists of 100M tokens extracted from Wikipedia articles with high rating. It features two different variants, namely, WikiText-2 and WikiText-103 which differ in the number of tokens and vocabulary size. WikiText-2 consists of 2M tokens and a vocabulary size of 33k whereas WikiText-103 is larger with 103M tokens and a vocabulary size of 267k.
  2. 2.
    LAMBADA Paperno et al. (2016): the LAMBADA dataset is extracted from the BookCorpus dataset boo and contains 10k passages. This dataset is useful for testing the ability of an LLM to capture long-range dependencies in text. The objective of this model is to predict the final word in a set of sentences, where humans need at least 50 tokens of context to accurately anticipate the word.
  3. 3.
    One Billion Word Benchmark Chelba et al. (2014) (1BW): this dataset contains one billion words extracted from the WMT 2011 News Crawl data and is used to measure progress in statistical language modeling.
  4. 4.
    WMT-14 French-English Translation Artetxe et al. (2018): This dataset contains 36 million training sentence pairs for english to french translation. The test set, which is used for evaluation purposes, consists of 3,003 sentence pairs.
  5. 5.
    Natural Questions Kwiatkowski et al. (2019): This dataset contains question-answer pairs from Google Search and Wikipedia-based annotations. The training, validation, and test sets consist of 307,372, 7,830, and 7,842 examples.
(Radford et al., 2019) demonstrates that these datasets are out-of-domain for the WebText dataset using Bloom filters containing 8-grams of the training tokens. While we have not conducted a similar study for the OpenWebText dataset, since both datasets are scraped and pre-processed using a similar method, we estimate that the percentages of 8 grams overlap with the training set will be similar.
Models: Table 6 describes the different GPT-2 configurations used in LFR.
124M 355M 774M 1.5B
Layers 12 24 36 48
#Heads 12 16 20 25
n_embd 768 1024 1280 1600
Table 6: Number of layers, attention heads, and the embedding dimensions in the GPT-2 models used for pretraining.
Pretraining: Table 7 shows the hyperparameters for pretraining the GPT-2 124M-1.5B parameter models.
Parameter Value
Context Length 1024
Embedding Dimension (768, 1024, 1280, 1600)
Total Iterations 40000
Effective Batch Size 512
Batch Size (16, 16, 8, 4)
Gradient Accumulation Steps (32, 32, 64, 128)
Block Size 1024
Weight Decay 1.00E-01
Adam $\beta_{1}$ 0.9
Adam $\beta_{2}$ 0.95
Warmup Iterations 2000
Minimum Learning Rate 6.00E-05
Maximum Learning Rate 6.00E-04
Learning Rate Schedule Linear
Learning Rate Decay Iterations 40000
GPUs (4xMI100, 4xMI210, 4xMI210, 4xMI210)
Table 7: Pretraining hyperparameters for the GPT-2 124M-1.5B parameter models. Parameters with multiple values (e.g. Embedding dimensions, batch size, gradient accumulation steps, and GPUs) specified in brackets are for the 124M, 345M, 774M, and 1.5B parameter models respectively.
Note that OpenAI pretrained the GPT-2 models using a batch size of 512. Due to insufficient GPU memory, we adjust the number of gradient accumulation steps to achieve the same effective batch size of 512.
Training Runtime Table 8 describes the runtime in days per pretraining experiment (40k iterations) for the different model configurations.
Model #Days GPUs
GPT-2 124M 2.2 4 AMD MI100
GPT-2 345M 8.8 4 AMD MI210
GPT-2 812M 14.1 4 AMD MI210
GPT-2 1.5B 20 4 AMD MI210
Table 8: Pretraining experiments runtime and GPU configurations used.
Finetuning: We use all the same hyperparameters as pretraining, except for the following:
  1. 1.
    Learning rate: 3.00E-5
  2. 2.
    Learning rate schedule: Constant
  3. 3.
    Total iterations: 50
A.2 Extended Analysis on Dropped and Retained Data Blocks
Model Mean Std Variance
GPT-2 124M 0.45 0.20 0.04
GPT-2 345M 0.30 0.12 0.01
Table 9: Mean, standard deviation (std), and variance of cosine similarity matrices for dropped data blocks across time scale (Phase 2 and Phase 4) for the GPT-2 124M and 345M models.
Model Mean Std Variance
GPT-2 124M 0.44 0.21 0.046
GPT-2 345M 0.32 0.13 0.018
Table 10: Mean, standard deviation (std), and variance of cosine similarity matrices for dropped and retained data blocks in Phase 4 of pretraining for the GPT-2 124M and 345M models.
Figure 4: Cosine similarity heatmaps for dropped and retained data blocks during Phase 4 of pretraining for the GPT-2 124M (right) and 345M (left) models.
Figure 5: Cosine similarity heatmap for dropped data blocks during Phase 4 of GPT-2 124M and Phase 2 of the 345M model.
Figure 6: Cosine similarity heatmaps for dropped and retained data blocks during Phase 2 of pretraining for the GPT-2 124M (right) and 345M (left) models.
Model Mean Std Variance
GPT-2 124M 0.42 0.19 0.04
GPT-2 345M 0.40 0.18 0.03
Table 11: Mean, standard deviation (std), and variance of cosine similarity matrices for dropped and retained data blocks in phase 2 of pretraining for the GPT-2 124M and 345M models.
Become a fan of Slate on Facebook. Follow us on Twitter.The first time I crocheted a soccer ball was on the occasion of the 2010 World Cup. It was being held on the continent of Africa, and I thought the African Flower hexagon motif was the perfect vehicle for a crochet soccer ball celebrating the continent’s first time hosting the World Cup: This time around, instead of using all 9000 of my favorite colors, I limited myself to the colors of the flags of the thirty-two countries that had made it to the final rounds of the World Cup competition, and I did my best to incorporate the designs of their flags into the thirty-two hexagons and pentagons of a soccer ball.
ML-77 Missile Launcher: Based on existing technology, the ML-77 is a rapid-fire missile launcher using seeking projectiles. Each projectile features a friend-or-foe recognition system, ensuring it will find a hostile target even if the user’s aim is not completely accurate. The locking mechanism of the ML-77 allows the shooter to ignore cover and line of sight when shooting at locked on enemies, though an attack roll is still required. Locking on to an enemy requires a move action when the enemy is in line of sight and lasts for the rest of the encounter, or until a new target is locked.
Table 12: Examples of text dropped by the 345M model in phase 2 (top) and phase 4 (bottom).
Unofficial reports claimed the car was powered by a 95kW 1.5-litre non-turbo petrol engine but Tada didn’t confirm. When asked what powers the S-FR Tada revealed he was considering three choices. "When you see the S-FR concept I suppose you imagine it is a 1.5-litre car but nowadays I can choose many kind of engines," he explained. "Downsized turbo, 1.5-litre naturally aspirated and something additional as well. Now we are thinking which one is the best engine for a small sports car." Tada also admitted that the company is unlikely to turn to a partner like it did with Subaru for the 86/BRZ or the new ’big brother’ sports car with BMW.
In April, MYIR released a Linux-powered MYS-6ULX single board computer, which was notable for being available in two different versions using NXP’s low power, Cortex-A7 i.MX6 UltraLite (UL) or the more affordable, and almost identical i.MX6 ULL SoC. Now, MYIR has released an “MYB-6ULX Expansion Board” designed to stack onto either model. The \$21.20 accessory adds a second 10100 Ethernet port to the MYS-6ULX, as well as new CAN, RS485, audio, micro-USB, RTC, and camera functions. MYB-6ULX Expansion Board with MYS-6ULX (left) and detail view (click images to enlarge). The MYB-6ULX Expansion Board has the same 70 x 55mm dimensions as the MYS-6ULX, which is available in two models: The i.MX6 UL based MYS-6ULX-IND has -40 to 85°C support instead of 0 to 70°C, and the i.MX6 ULL based MYS-6ULX-IOT features a USB-powered WiFi radio. The 4-layer expansion board runs on 5V power, and shares the industrial temperature support of the IND model.
Table 13: Examples of text retained by the 345M model in Phase 2 (top) and Phase 4 (bottom).
In the book, the mythical California is ruled by Queen Califa and populated only with female warriors who brandish gold weapons. They even harness their animals in gold because it is the only mineral on the island. The legend of Califa and her island was well known among New World explorers. In 1536 when Hernán Cortéz arrived in Baja California, he believed he had landed on the legendary island. Over three hundred years later gold was discovered in California, making the legend partially true and earning the state its nickname: The Golden State.
Segregated Witness, defined by Bitcoin Improvement Proposal 141 (BIP141), was deployed using an activation mechanism (BIP9) that requires 95 percent of all miners (by hash power) to signal support for the upgrade within the span of a two-week difficulty period. That’s at least 1916 blocks within 2016 blocks, to be exact. This threshold has just been reached. While the current difficulty period will not end until tomorrow, all blocks in this difficulty period are signaling support for the upgrade so far. This now totals over 1916 of them.
Table 14: Examples of text dropped by the 124M model in Phase 2 (top) and Phase 4 (bottom).
to the GUI installer. Most notably there is no support for configuring partition layout, storage methods or package selection. Please refer to the official documentation for details. Here you can find some useful information on creating and using kickstart files which can be used to perform advanced configuring without the need for the GUI installer. The message "Insufficient memory to configure kdump!" appears during install. This is a known issue which appears on systems with less than 2 GB RAM. This can be ignored. Content for both the i386 and x86_64 architectures is split into two DVDs. We have tried to get all basic server and basic desktop installs only from DVD-1. Make sure that you setup correctly the selinux context of the public key if you transfer it to a CentOS 6 server with selinux enabled.
Once you signed up, you can either click on the Todo tab or the sign in link to gain access to the application. Note that if you are selecting sign in in the same session in which you signed up, you will automatically sign in with the same account you used for signing up. If you are signing in during a new session, you will be presented with Azure AD’s credentials prompt: sign in using an account compatible with the sign up option you chose earlier (the exact same account if you used user consent, any user form the same tenant if you used admin consent). If you try to sign-in before the tenant administrator has provisioned the app in the tenant using the Sign up link above, you will see the following error.
Table 15: Examples of text retained by the 124M model in phase 2 (top) and phase 4 (bottom).
In this section, we expand on the ablation study in Section 5.8 in order to better characterize the data blocks considered easy / hard.
Tables 12 and 14 provides examples of text blocks dropped in Phases 2 and 4 by the 345M and 124M parameter models respectively. Similarly, Tables 13 and 15 provide examples of data blocks retained by the models in Phases 2 and 4. We printed out and went over all the text dropped and retained in both Phases, and notice that text considered easy in phase 2 was more conversational, and those considered easy in phase 4 were more factual. This might indicate that the model first learned conversations and personal anecdotes, before being able to retain factual information. These findings are further corroborated by the examples of data retained in both phases. We are working on further analysis across different model families and sizes to strengthen this understanding.
Next, we continue the analysis of the cosine similarity heatmaps evaluated across training time and model parameter scales presented in Section 5.8. Here, we answer the following questions:
  1. 1.
    Are there similarities in the data blocks considered easy and dropped in Phase 4 of training of the 124M parameter model with those considered easy and dropped by the 345M parameter model in Phase 2?
  2. 2.
    Are the data blocks dropped similar to those retained at any given training phase? Note that Section 5.8 presented this analysis only for Phase 4 of the 124M and 345M parameter models in Figure 4.
Figure 5 depicts the cosine similarity heatmap for the data blocks dropped by the 124M parameter model in Phase 4 (X-axis) with those dropped by the 345M parameter model in Phase 2 (Y-axis). The mean, standard deviation, and variance of the similarity are 0.43, 0.18, and 0.03 respectively. In contrast, the mean cosine similarity of data blocks dropped in Phase 2 of pretraining of both the models was 0.38 (Section 5.8 and Figure 3). This indicates that the smaller model "catches up" with the knowledge accumulated by the larger model, and considers similar block easy in Phase 4 as those considered easy by the larger model in Phase 2.
Next, in Figure 6, we plot the cosine similarity heatmap for the dropped and retained data blocks in Phase 2 for the 124M (left) and 345M (right) parameter models. The mean, variance, and standard deviations of the similarity are shown in Table 11. Observing the mean similarity value and heatmap in Table 10 and Figure 4, we find that the cosine similarity for dropped and retained data blocks is higher in Phase 2 than Phase 4 in case of the 345M parameter model. In contrast, the value remains high in both Phases for the 124M parameter model. This finding indicates that both the smaller and larger model start the training by being confused about similar data blocks. However, the larger capacity of the 345M parameter model allows it to learn the dataset well in Phases 2 and 3, and move on to more complex data blocks in Phase 4 (thus reducing the mean similarity in Phase 4). The smaller model continues remaining unsure about similar data blocks. Since we observed that the smaller model "catches up" with the training of the larger model (in Figure 5), we hypothesize that the smaller model will eventually display similar behaviour as the larger model once trained for longer iterations.